# Electrically Tunable and Enhanced Nonlinearity of Moiré Exciton-Polaritons in Transition Metal Dichalcogenide Bilayers

### v2

@@ Acknowledgements.
Acknowledgments.—We acknowledge the support from UK EPSRC Awards No. EP/X017222/1.
@@
Appendix A Interlayer and intralayer exciton transition matrix element
@@
$\mathcal{X}^{\dagger}_{\sigma}=\sum_{s\alpha}\sum_{\mathbf{Q}}C^{s}_{\alpha\sigma}(\mathbf{Q})X^{s,\dagger}_{\alpha\sigma}(\mathbf{Q}).$ (12)
@@
$(E^{s}_{\alpha\sigma}(\mathbf{Q})-\mathcal{E}_{\sigma})C^{s}_{\alpha\sigma}(\mathbf{Q})\!=\!\!\!\sum_{s^{\prime}\alpha^{\prime}\mathbf{Q}^{\prime}}\!\!\!W^{s^{\prime}s}_{\alpha^{\prime}\alpha}(\mathbf{Q}^{\prime},\mathbf{Q})C^{s^{\prime}}_{\alpha^{\prime}\sigma}(\mathbf{Q}^{\prime})$ (13)
@@
$W^{s^{\prime}s}_{\alpha^{\prime}\alpha}(\mathbf{Q}^{\prime},\mathbf{Q})=\langle 0|X^{s^{\prime}}_{\alpha^{\prime}\sigma}(\mathbf{Q}^{\prime})\mathcal{H}_{t}X^{s\dagger}_{\alpha\sigma}(\mathbf{Q})|0\rangle.$ (14)
This determines how the interlayer exciton and intralayer exciton are hybridized. Solving Eq. (13), we obtain the hybridized exciton in Eq. (12) with energy $\mathcal{E}_{\sigma}$.
@@
$-\sum_{\mathbf{k}}\bar{\phi}_{\alpha\sigma}^{s^{\prime}}(\mathbf{k}-\zeta^{c}_{s^{\prime}}\mathbf{Q}^{\prime})T^{v}_{\ell_{v}^{\prime}\ell_{v}}(\mathbf{k}-\mathbf{Q}+\delta\mathbf{K},\mathbf{k}-\mathbf{Q}^{\prime}+\delta\mathbf{K})\delta_{\ell_{c}^{\prime}\ell_{c}}\phi_{\alpha\sigma}^{s}(\mathbf{k}-\zeta^{c}_{s}\mathbf{Q}),$ (15)
@@
To construct the hybridized exciton equation in Eq. (13), we approximate hopping constant [51, 50] as in Eq. (5)
@@
$\mathbf{G}_{\ell}=\mathbf{G}_{\ell}^{(\eta)}=(-1)^{\eta}(C^{\eta}_{3}-C^{0}_{3})\mathbf{K}_{\ell},\quad(\eta=1,2,3)$ (17)
@@
$t^{v}(\mathbf{K}_{\ell}+\mathbf{G}^{(\eta)}_{\ell})=$ $t^{v}\begin{cases}1,&\text{(P stacking)}\\ \mathrm{e}^{-i\frac{4}{3}\pi\eta},&\text{(AP stacking)}\end{cases}$ (18)
@@
$T^{c}_{\ell^{\prime}\ell}(\mathbf{k}^{\prime},\mathbf{k})=t^{c}\sum_{\eta=0}^{2}\delta_{\mathbf{k}^{\prime}-\mathbf{k},\mathbf{b}_{\eta}},\quad T^{v}_{\ell^{\prime}\ell}(\mathbf{k}^{\prime},\mathbf{k})=t^{v}\sum_{\eta=0}^{2}\delta_{\mathbf{k}^{\prime}-\mathbf{k},\mathbf{b}_{\eta}}\begin{cases}1,&\text{(P stacking)}\\ \mathrm{e}^{-i\frac{4}{3}\pi\eta},&\text{(AP stacking)}\end{cases}$ (19)
with $\mathbf{b}_{0}=0$. We note that this result has been obtained in Ref. [50]. In their notation, the momentum $\mathbf{k}$ and $\mathbf{k}^{\prime}$ are measured from the valleys $\mathbf{K}_{1,2}$, while in our paper we measured the momentum from the $\Gamma$ point. With this hopping constant, this reduces Eq. (A) (P stacking) to
@@
$-t_{v}\bar{\phi}_{\alpha^{\prime}\sigma}^{s^{\prime}}[\mathbf{k}\!-\!\zeta^{c}_{s^{\prime}}(\bar{\mathbf{Q}}^{\prime}\!+\!\bm{g}^{\prime})]\phi_{\alpha\sigma}^{s}[\mathbf{k}\!-\!\zeta^{c}_{s}(\bar{\mathbf{Q}}\!+\!\bm{g})]\delta_{\ell_{c}\ell_{c}^{\prime}}\tau^{x}_{\ell_{v}^{\prime}\ell_{v}}\Big{]}\sum_{\eta=0}^{2}\delta_{\bm{g}^{\prime}-\bm{g},\mathbf{b}_{\eta}-(\bar{\mathbf{Q}}^{\prime}-\bar{\mathbf{Q}})},$ (20)
where $\mathbf{Q}=\bar{\mathbf{Q}}+\bm{g}$ with $\bar{\mathbf{Q}}$ being the momentum restricted in the mini Brillouin zone of the moiré superlattice (see Fig. 2a). To further simplify the above, we observe that $\bm{g}=i\mathbf{b}_{1}+j\mathbf{b}_{2}$ is also reciprocal lattice vector of a moiré lattice. This implies that Eq. (A) is zero if $\bar{\mathbf{Q}}^{\prime}\neq\bar{\mathbf{Q}}$. This gives
$W^{s^{\prime}s}_{\alpha^{\prime}\alpha}(\mathbf{Q},\mathbf{Q}^{\prime})=w_{\nu^{\prime}\nu}(\bar{\mathbf{Q}})\delta_{\bar{\mathbf{Q}}^{\prime},\bar{\mathbf{Q}}}.$ (21)
@@
$w_{\nu^{\prime}\nu}(\bar{\mathbf{Q}})=$ $(t_{c}M^{v}_{\nu^{\prime}\nu}(\bar{\mathbf{Q}})\delta_{\ell_{v}^{\prime}\ell_{v}}\tau^{x}_{\ell_{c}^{\prime}\ell_{c}}-t_{v}M^{c}_{\nu^{\prime}\nu}(\bar{\mathbf{Q}})\delta_{\ell_{c}^{\prime}\ell_{c}}\tau^{x}_{\ell_{v}^{\prime}\ell_{v}})\sum_{\eta=0}^{2}\delta_{\bm{g}^{\prime}-\bm{g},\mathbf{b}_{\eta}}.$ (22)
@@
The summation of the delta functions in Eq. (22) correspond to the mixing of the seven mBZs shown in the inset of Fig. 2a. One may include more mixing beyond these mBZ. However, this involves the electrons with higher-energy dispersion and their effects will be suppressed.
Due to the conservation of $\bar{\mathbf{Q}}$ in Eq. (21), each different $\bar{\mathbf{Q}}$ modes in Eq. (13) do not mix together and break into independent sectors. This yields Eq. (7),
$[E_{\nu\sigma}(\bar{\mathbf{Q}})-\mathcal{E}^{\bar{\alpha}}_{\sigma}(\bar{\mathbf{Q}})]C^{\bar{\alpha}}_{\nu\sigma}(\bar{\mathbf{Q}})\!=\!\sum_{\nu^{\prime}}w_{\nu^{\prime}\nu}(\bar{\mathbf{Q}})C^{\bar{\alpha}}_{\nu^{\prime}\sigma}(\bar{\mathbf{Q}}),$ (23)
@@ Appendix B X-X interaction
$\Omega^{\bar{\beta}\bar{\alpha}}_{\sigma}=\langle 0|\mathcal{X}^{\bar{\alpha}}_{\sigma}\mathcal{X}^{\bar{\beta}}_{\sigma}\mathcal{H}\mathcal{X}^{\bar{\beta},\dagger}_{\sigma}\mathcal{X}^{\bar{\alpha},\dagger}_{\sigma}|0\rangle.$ (24)
@@
$\sum_{\alpha\mathbf{Q}}\bar{\Phi}^{\ell_{c}\ell_{v}}_{\alpha\mathbf{Q},\sigma}(\mathbf{k}_{c}^{\prime},\mathbf{k}_{v}^{\prime})\Phi^{\ell_{c}\ell_{v}}_{\alpha\mathbf{Q},\sigma}(\mathbf{k}_{c},\mathbf{k}_{v})=\delta_{\mathbf{k}_{c}\mathbf{k}_{c}^{\prime}}\delta_{\mathbf{k}_{v}\mathbf{k}_{v}^{\prime}},\quad\sum_{\alpha\mathbf{Q}}\bar{\Phi}^{\ell_{c}\ell_{v}}_{\alpha\mathbf{Q},\sigma}(\mathbf{k}_{c},\mathbf{k}_{v})X^{\ell_{c}\ell_{v}}_{\alpha\mathbf{Q},\sigma}=a^{\ell_{c}\dagger}_{\mathbf{k}_{c}\sigma}b^{\ell_{v}}_{\mathbf{k}_{v}\sigma}.$ (25)
@@
$\Omega^{mn}_{\sigma}=\mathcal{E}^{\bar{\beta}}_{\sigma}+\mathcal{E}^{\bar{\alpha}}_{\sigma}+\Delta^{mn}_{\sigma},$ (26)
@@
$[X^{s^{\prime}}_{\alpha^{\prime}\sigma}(\mathbf{Q}^{\prime}),X^{s,\dagger}_{\alpha\sigma}(\mathbf{Q})]=\delta_{\alpha\alpha^{\prime}}\delta_{ss^{\prime}}\delta_{\mathbf{Q},\mathbf{Q}^{\prime}}-D^{s^{\prime}s}_{\alpha^{\prime}\alpha}(\mathbf{Q}^{\prime},\mathbf{Q})$ (27)
@@
$D^{s^{\prime}s}_{\alpha^{\prime}\alpha}(\mathbf{Q}^{\prime},\mathbf{Q})=$ $\sum_{\mathbf{k}_{c}\mathbf{k}_{c}^{\prime}}\sum_{\mathbf{k}_{v}\mathbf{k}_{v}^{\prime}}\Phi^{s}_{\alpha\mathbf{Q}}(\mathbf{k}_{c},\mathbf{k}_{v})\bar{\Phi}^{s^{\prime}}_{\alpha^{\prime}\mathbf{Q}^{\prime}}(\mathbf{k}_{c}^{\prime},\mathbf{k}_{v}^{\prime})\Big{\{}\delta_{\ell_{v}\ell_{v}^{\prime}}\delta_{\mathbf{k}_{v}\mathbf{k}_{v}^{\prime}}a^{\ell_{c}\dagger}_{\mathbf{k}_{c}}a^{\ell_{c}^{\prime}}_{\mathbf{k}_{c}^{\prime}}+\delta_{\ell_{c}\ell_{c}^{\prime}}\delta_{\mathbf{k}_{c}^{\prime}\mathbf{k}_{c}}b^{\ell_{v}}_{\mathbf{k}_{v}}b^{\ell_{v}^{\prime}\dagger}_{\mathbf{k}_{v}^{\prime}}\Big{\}}.$ (28)
@@
$-\langle 0|X^{\tilde{s}^{\prime}}_{\tilde{\alpha}^{\prime}\sigma}(\tilde{\mathbf{Q}}^{\prime})D^{s^{\prime}s}_{\alpha^{\prime}\alpha}(\mathbf{Q}^{\prime},\mathbf{Q})X^{\tilde{s},\dagger}_{\tilde{\alpha}\sigma}(\tilde{\mathbf{Q}})|0\rangle.$ (29)
Using $D^{s^{\prime}s}_{\alpha^{\prime}\alpha}(\mathbf{Q}^{\prime},\mathbf{Q})|0\rangle=0$, the commutation relation Eq. (27), and
@@
$+\delta_{\ell_{c}\ell_{c}^{\prime}}\delta_{\ell_{v}^{\prime}\tilde{\ell}_{v}}\bar{\Phi}^{s^{\prime}}_{\alpha^{\prime}\mathbf{Q}^{\prime}\sigma}(\mathbf{k}_{c},\tilde{\mathbf{k}}_{v})\Phi^{\tilde{\ell}_{c}\ell_{v}}_{\beta^{\prime},\tilde{\mathbf{k}}_{c}-\mathbf{k}_{v}}(\tilde{\mathbf{k}}_{c},\mathbf{k}_{v})X^{\tilde{\ell}_{c}\ell_{v},\dagger}_{\beta^{\prime},\tilde{\mathbf{k}}_{c}-\mathbf{k}_{v}}\Big{\}}.$ (30)
@@
Appendix C Momentum integration in direct and exchange interactions

### v3

@@ Acknowledgements.
Acknowledgments.—We acknowledge the support from UK EPSRC Awards No. EP/X017222/1. K. W. S. is supported by Xiamen University Malaysia Research Fund (Grant No. XMUMRF/2025-C15/IPHY/0005).
@@
I Interlayer and intralayer exciton transition matrix element
@@
$\mathcal{X}^{\dagger}_{\sigma}=\sum_{s\alpha}\sum_{\mathbf{Q}}C^{s}_{\alpha\sigma}(\mathbf{Q})X^{s,\dagger}_{\alpha\sigma}(\mathbf{Q}).$ (S1)
@@
$(E^{s}_{\alpha\sigma}(\mathbf{Q})-\mathcal{E}_{\sigma})C^{s}_{\alpha\sigma}(\mathbf{Q})\!=\!\!\!\sum_{s^{\prime}\alpha^{\prime}\mathbf{Q}^{\prime}}\!\!\!W^{s^{\prime}s}_{\alpha^{\prime}\alpha}(\mathbf{Q}^{\prime},\mathbf{Q})C^{s^{\prime}}_{\alpha^{\prime}\sigma}(\mathbf{Q}^{\prime})$ (S2)
@@
$W^{s^{\prime}s}_{\alpha^{\prime}\alpha}(\mathbf{Q}^{\prime},\mathbf{Q})=\langle 0|X^{s^{\prime}}_{\alpha^{\prime}\sigma}(\mathbf{Q}^{\prime})\mathcal{H}_{t}X^{s\dagger}_{\alpha\sigma}(\mathbf{Q})|0\rangle.$ (S3)
This determines how the interlayer exciton and intralayer exciton are hybridized. Solving Eq. (S2), we obtain the hybridized exciton in Eq. (S1) with energy $\mathcal{E}_{\sigma}$.
@@
$-\sum_{\mathbf{k}}\bar{\phi}_{\alpha\sigma}^{s^{\prime}}(\mathbf{k}-\zeta^{c}_{s^{\prime}}\mathbf{Q}^{\prime})T^{v}_{\ell_{v}^{\prime}\ell_{v}}(\mathbf{k}-\mathbf{Q}+\delta\mathbf{K},\mathbf{k}-\mathbf{Q}^{\prime}+\delta\mathbf{K})\delta_{\ell_{c}^{\prime}\ell_{c}}\phi_{\alpha\sigma}^{s}(\mathbf{k}-\zeta^{c}_{s}\mathbf{Q}),$ (S4)
@@
To construct the hybridized exciton equation in Eq. (S2), we approximate hopping constant [52, 51] as in Eq. (5)
$T^{c,v}_{\ell^{\prime}\ell}(\mathbf{k}^{\prime},\mathbf{k})=$ $\sum_{\mathbf{G}_{\ell}\mathbf{G}_{\ell^{\prime}}}t^{c,v}(\mathbf{K}_{\ell}+\mathbf{G}_{\ell})\delta_{\mathbf{k}^{\prime}-\mathbf{k},\mathbf{G}_{\ell^{\prime}}-\mathbf{G}_{\ell}}\tau^{x}_{\ell\ell^{\prime}},$ (S5)
@@
$\mathbf{G}_{\ell}=\mathbf{G}_{\ell}^{(\eta)}=(-1)^{\eta}(C^{\eta}_{3}-C^{0}_{3})\mathbf{K}_{\ell},\quad(\eta=1,2,3)$ (S6)
@@
$t^{v}(\mathbf{K}_{\ell}+\mathbf{G}^{(\eta)}_{\ell})=$ $t^{v}\begin{cases}1,&\text{(P stacking)}\\ \mathrm{e}^{-i\frac{4}{3}\pi\eta},&\text{(AP stacking)}\end{cases}$ (S7)
@@
$T^{c}_{\ell^{\prime}\ell}(\mathbf{k}^{\prime},\mathbf{k})=t^{c}\sum_{\eta=0}^{2}\delta_{\mathbf{k}^{\prime}-\mathbf{k},\mathbf{b}_{\eta}},\quad T^{v}_{\ell^{\prime}\ell}(\mathbf{k}^{\prime},\mathbf{k})=t^{v}\sum_{\eta=0}^{2}\delta_{\mathbf{k}^{\prime}-\mathbf{k},\mathbf{b}_{\eta}}\begin{cases}1,&\text{(P stacking)}\\ \mathrm{e}^{-i\frac{4}{3}\pi\eta},&\text{(AP stacking)}\end{cases}$ (S8)
with $\mathbf{b}_{0}=0$. We note that this result has been obtained in Ref. [51]. In their notation, the momentum $\mathbf{k}$ and $\mathbf{k}^{\prime}$ are measured from the valleys $\mathbf{K}_{1,2}$, while in our paper we measured the momentum from the $\Gamma$ point. With this hopping constant, this reduces Eq. (I) (P stacking) to
@@
$-t_{v}\bar{\phi}_{\alpha^{\prime}\sigma}^{s^{\prime}}[\mathbf{k}\!-\!\zeta^{c}_{s^{\prime}}(\bar{\mathbf{Q}}^{\prime}\!+\!\bm{g}^{\prime})]\phi_{\alpha\sigma}^{s}[\mathbf{k}\!-\!\zeta^{c}_{s}(\bar{\mathbf{Q}}\!+\!\bm{g})]\delta_{\ell_{c}\ell_{c}^{\prime}}\tau^{x}_{\ell_{v}^{\prime}\ell_{v}}\Big{]}\sum_{\eta=0}^{2}\delta_{\bm{g}^{\prime}-\bm{g},\mathbf{b}_{\eta}-(\bar{\mathbf{Q}}^{\prime}-\bar{\mathbf{Q}})},$ (S9)
where $\mathbf{Q}=\bar{\mathbf{Q}}+\bm{g}$ with $\bar{\mathbf{Q}}$ being the momentum restricted in the mini Brillouin zone of the moiré superlattice (see Fig. 2a). To further simplify the above, we observe that $\bm{g}=i\mathbf{b}_{1}+j\mathbf{b}_{2}$ is also reciprocal lattice vector of a moiré lattice. This implies that Eq. (I) is zero if $\bar{\mathbf{Q}}^{\prime}\neq\bar{\mathbf{Q}}$. This gives
$W^{s^{\prime}s}_{\alpha^{\prime}\alpha}(\mathbf{Q},\mathbf{Q}^{\prime})=w_{\nu^{\prime}\nu}(\bar{\mathbf{Q}})\delta_{\bar{\mathbf{Q}}^{\prime},\bar{\mathbf{Q}}}.$ (S10)
@@
$w_{\nu^{\prime}\nu}(\bar{\mathbf{Q}})=$ $(t_{c}M^{v}_{\nu^{\prime}\nu}(\bar{\mathbf{Q}})\delta_{\ell_{v}^{\prime}\ell_{v}}\tau^{x}_{\ell_{c}^{\prime}\ell_{c}}-t_{v}M^{c}_{\nu^{\prime}\nu}(\bar{\mathbf{Q}})\delta_{\ell_{c}^{\prime}\ell_{c}}\tau^{x}_{\ell_{v}^{\prime}\ell_{v}})\sum_{\eta=0}^{2}\delta_{\bm{g}^{\prime}-\bm{g},\mathbf{b}_{\eta}}.$ (S11)
@@
The summation of the delta functions in Eq. (S11) correspond to the mixing of the seven mBZs shown in the inset of Fig. 2a. One may include more mixing beyond these mBZ. However, this involves the electrons with higher-energy dispersion and their effects will be suppressed.
Due to the conservation of $\bar{\mathbf{Q}}$ in Eq. (S10), each different $\bar{\mathbf{Q}}$ modes in Eq. (S2) do not mix together and break into independent sectors. This yields Eq. (7),
$[E_{\nu\sigma}(\bar{\mathbf{Q}})-\mathcal{E}^{\bar{\alpha}}_{\sigma}(\bar{\mathbf{Q}})]C^{\bar{\alpha}}_{\nu\sigma}(\bar{\mathbf{Q}})\!=\!\sum_{\nu^{\prime}}w_{\nu^{\prime}\nu}(\bar{\mathbf{Q}})C^{\bar{\alpha}}_{\nu^{\prime}\sigma}(\bar{\mathbf{Q}}),$ (S12)
@@ II X-X interaction
$\Omega^{\bar{\beta}\bar{\alpha}}_{\sigma}=\langle 0|\mathcal{X}^{\bar{\alpha}}_{\sigma}\mathcal{X}^{\bar{\beta}}_{\sigma}\mathcal{H}\mathcal{X}^{\bar{\beta},\dagger}_{\sigma}\mathcal{X}^{\bar{\alpha},\dagger}_{\sigma}|0\rangle.$ (S13)
@@
$\sum_{\alpha\mathbf{Q}}\bar{\Phi}^{\ell_{c}\ell_{v}}_{\alpha\mathbf{Q},\sigma}(\mathbf{k}_{c}^{\prime},\mathbf{k}_{v}^{\prime})\Phi^{\ell_{c}\ell_{v}}_{\alpha\mathbf{Q},\sigma}(\mathbf{k}_{c},\mathbf{k}_{v})=\delta_{\mathbf{k}_{c}\mathbf{k}_{c}^{\prime}}\delta_{\mathbf{k}_{v}\mathbf{k}_{v}^{\prime}},\quad\sum_{\alpha\mathbf{Q}}\bar{\Phi}^{\ell_{c}\ell_{v}}_{\alpha\mathbf{Q},\sigma}(\mathbf{k}_{c},\mathbf{k}_{v})X^{\ell_{c}\ell_{v}}_{\alpha\mathbf{Q},\sigma}=a^{\ell_{c}\dagger}_{\mathbf{k}_{c}\sigma}b^{\ell_{v}}_{\mathbf{k}_{v}\sigma}.$ (S14)
@@
$\Omega^{mn}_{\sigma}=\mathcal{E}^{\bar{\beta}}_{\sigma}+\mathcal{E}^{\bar{\alpha}}_{\sigma}+\Delta^{mn}_{\sigma},$ (S15)
@@
$[X^{s^{\prime}}_{\alpha^{\prime}\sigma}(\mathbf{Q}^{\prime}),X^{s,\dagger}_{\alpha\sigma}(\mathbf{Q})]=\delta_{\alpha\alpha^{\prime}}\delta_{ss^{\prime}}\delta_{\mathbf{Q},\mathbf{Q}^{\prime}}-D^{s^{\prime}s}_{\alpha^{\prime}\alpha}(\mathbf{Q}^{\prime},\mathbf{Q})$ (S16)
@@
$D^{s^{\prime}s}_{\alpha^{\prime}\alpha}(\mathbf{Q}^{\prime},\mathbf{Q})=$ $\sum_{\mathbf{k}_{c}\mathbf{k}_{c}^{\prime}}\sum_{\mathbf{k}_{v}\mathbf{k}_{v}^{\prime}}\Phi^{s}_{\alpha\mathbf{Q}}(\mathbf{k}_{c},\mathbf{k}_{v})\bar{\Phi}^{s^{\prime}}_{\alpha^{\prime}\mathbf{Q}^{\prime}}(\mathbf{k}_{c}^{\prime},\mathbf{k}_{v}^{\prime})\Big{\{}\delta_{\ell_{v}\ell_{v}^{\prime}}\delta_{\mathbf{k}_{v}\mathbf{k}_{v}^{\prime}}a^{\ell_{c}\dagger}_{\mathbf{k}_{c}}a^{\ell_{c}^{\prime}}_{\mathbf{k}_{c}^{\prime}}+\delta_{\ell_{c}\ell_{c}^{\prime}}\delta_{\mathbf{k}_{c}^{\prime}\mathbf{k}_{c}}b^{\ell_{v}}_{\mathbf{k}_{v}}b^{\ell_{v}^{\prime}\dagger}_{\mathbf{k}_{v}^{\prime}}\Big{\}}.$ (S17)
@@
$-\langle 0|X^{\tilde{s}^{\prime}}_{\tilde{\alpha}^{\prime}\sigma}(\tilde{\mathbf{Q}}^{\prime})D^{s^{\prime}s}_{\alpha^{\prime}\alpha}(\mathbf{Q}^{\prime},\mathbf{Q})X^{\tilde{s},\dagger}_{\tilde{\alpha}\sigma}(\tilde{\mathbf{Q}})|0\rangle.$ (S18)
Using $D^{s^{\prime}s}_{\alpha^{\prime}\alpha}(\mathbf{Q}^{\prime},\mathbf{Q})|0\rangle=0$, the commutation relation Eq. (S16), and
@@
$+\delta_{\ell_{c}\ell_{c}^{\prime}}\delta_{\ell_{v}^{\prime}\tilde{\ell}_{v}}\bar{\Phi}^{s^{\prime}}_{\alpha^{\prime}\mathbf{Q}^{\prime}\sigma}(\mathbf{k}_{c},\tilde{\mathbf{k}}_{v})\Phi^{\tilde{\ell}_{c}\ell_{v}}_{\beta^{\prime},\tilde{\mathbf{k}}_{c}-\mathbf{k}_{v}}(\tilde{\mathbf{k}}_{c},\mathbf{k}_{v})X^{\tilde{\ell}_{c}\ell_{v},\dagger}_{\beta^{\prime},\tilde{\mathbf{k}}_{c}-\mathbf{k}_{v}}\Big{\}}.$ (S19)
@@
III Momentum integration in direct and exchange interactions